OpenStudy (anonymous):

One solution of the equation 3x2-16x+5=0 is 1/3. What is the other solution?

OpenStudy (anonymous):

factors as \[(3x-1)(something)\] figure out the \(something\)

OpenStudy (mathmale):

If you know synthetic division, please use 1/3 as the divisor and try synth. div. Otherwise, convert that 1/3 to a factor: (3x-1); divide the polynomial by that.

OpenStudy (anonymous):

or just figure it out by thinking \[(3x-1)(ax+b)=3x^2-16x+5\] pretty clear that it has to by \(x\) for the first term and \(-5\) for the second since \(3x\times x=3x^2\) and \(-1\times -5=5\)

OpenStudy (anonymous):

O ok I get it thanks